Which of the following equations have the same solution? Solve all three equations separately and explain each step.
I. 4x – 14 = 2x + 8
II. 3x – 21 = 12
III. x-2/5=4x+2/25
I just need help on the third equation the first two both equal 11

Answer by ewatrrr(24785)

 
Hi,
III. x-2/5=4x+2/25
-2/5 - 2/25 = 3x
-10/25 - 2/25 = 3x
-12/25 = 3x
-12/75 = x